A young star with a luminosity 150L⊙ outflow velocity of 30 km s−1, and a mass loss rate of M ̇ = 10−4M⊙ yr−1 Does radiation pressure play a role in driving the outflow?



1
Expert's answer
2021-10-31T16:36:39-0400

"p=\frac{\sigma v^2 M}{cL}=3.6\cdot 10^{-21}~Pa,"

the radiation pressure does not play a role in driving the outflow.